How many ounces of a 22% alcohol solution mixed with how many ounces of pure alcohol would form 60 ounces of a 35% solution?
Answer by Paul(988)

0.22(x+60)=0.35(60)
0.22x+13.2=21
0.22x=7.8
x=35.45
Hence, about 35.45 ounces of alcohol should be mixed.
PAul.